OpenStudy (anonymous):

Shannon makes large cookies. These cookies are 22 inches in diameter. How much icing will she need to cover the top of each cookie? Answer 11π in2 22π in2 484π in2 121π in2

OpenStudy (across):

\[A=\pi r^2\]

OpenStudy (across):

\[r=\frac{d}{2}\]

OpenStudy (apoorvk):

radius of each cookie = diameter/2 = 11 inches (those are pretty big cookies by the way!!) so area of the surface of each cookie = pi times (radius)^2 so area = 11*11*22/7 = 38.14 square inches.